OpenStudy (anonymous):

What is the percent by mass of water in MgSO4 • 7H2O? A. 51.1% B. 195% C. 56.0% D. 21.0%

OpenStudy (anonymous):

M = molar mass(gmol^-1) M(MgSO4.7H2O) = [24+32+(4x16)+(1x14)+(7x16)]gmol^-1 = 246gmol^-1 M(H2O) = 7X(16+2) = 126gmol^-1 %H2O = (126/246)X100 = 51.2195122% aNSWER = 51.2%(3s.f.) Answer Is A